Determine the volume in liters occupied by 14g of nitrogen gas at STP?
1
Expert's answer
2020-01-20T05:10:45-0500

pV=nRT, V= nRT/p.

n=m/Mr.

n=14/28=0.5 mol.

V=0.5 x 8.314 x 273/101.325= 11.2 L.